# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v2

@@ Abstract
Convex polytopes have compact representations and exhibit convexity, which makes them suitable for abstracting obstacle-free spaces from various environments. Existing methods for generating convex polytopes always struggle to strike a balance between two requirements, producing high-quality polytope and efficiency. Moreover, another crucial requirement for convex polytopes to accurately contain certain seed point sets, such as a robot or a front-end path, is proposed in various tasks, which we refer to as manageability. In this paper, we show that we can achieve generation of high-quality convex polytope while ensuring both efficiency and manageability simultaneously, by introducing Fast Iterative Regional Inflation (FIRI). FIRI consists of two iteratively executed submodules: Restrictive Inflation (RsI) and computation of the Maximum Volume Inscribed Ellipsoid (MVIE) of convex polytope. By explicitly incorporating constraints that include the seed point set, RsI guarantees manageability. Meanwhile, the iterative monotonic optimization of MVIE, which serves as a lower bound of the volume of convex polytope, ensures high-quality results of FIRI. In terms of efficiency, we design methods tailored to the low-dimensional and multi-constrained nature of both modules, resulting in orders of magnitude improvement compared to generic solvers. Notably, for 2-D MVIE, we present a novel analytical algorithm that achieves linear-time complexity for the first time, further enhancing the efficiency of FIRI in the 2-D scenario. Extensive benchmarks conducted against state-of-the-art methods validate the superior performance of FIRI in terms of quality, manageability, and efficiency. Furthermore, various real-world applications showcase the generality and practicality of FIRI. The high-performance code of FIRI will be open-sourced for the reference of the community.
@@ I Introduction
In robotics, a key task is to navigate without collisions, which involves frequent interactions with environments abundant in vast amounts of discrete obstacle information. For this interaction requirement, the use of convex polytopes is appealing because they offer a compact and structured geometric abstraction of the feasible space, greatly alleviating the burden of collision avoidance [1, 2, 3]. Furthermore, this abstraction is advantageous for storing large-scale environmental feasible spaces and enables further analysis of topological information, such as the construction of road maps [4, 5]. Additionally, convex polytopes construct convex linear safety constraints from non-convex obstacles, which benefits problem formulation and solution. In fact, in several applications, this approach even transforms the problem into convex optimization, leading to the attainment of global optima [6, 7].
@@
Although convex polytope provides a compact and simple representation, generating satisfactory convex polytopes is far from a trivial task. First, a larger convex polytope implies a more comprehensive extraction of information from the safety space, which is advantageous for tasks requiring the spatial search within the safety space, such as trajectory planning. As illustrated in Fig. 1(a), a larger convex polytope facilitates the generation of a smoother trajectory. Therefore, we aim for the generated polytope to have the largest possible volume, and refer to the volume of the polytope as its quality in this paper. Furthermore, maximizing computational efficiency extends the application of convex polytope to online high-speed tasks or limited onboard computational resources. Despite extensive research  [1, 8, 9, 2] efforts dedicated to computing free convex hull from environments, striking a trade-off between the quality of the convex polytope and the efficiency of the generation process remains a formidable challenge. Typically, either existing approaches excel at generating high-quality free convex regions but require substantial computational budget [1, 8], or they demand a cheap computational overhead yet yield conservative results [9, 2].
@@
In addition to quality and efficiency, several applications impose a crucial requirement for the generated convex polytope to accurately encompass a specified set of points, to which we refer as the manageability. For example, in the trajectory planning depicted in Fig. 1(b), the corridor-based approach [2, 10] typically involves using the path composed of multiple blue line segments generated by path planning as the seeds to generate multiple convex hulls, which collectively form a safety corridor. The trajectory is then constrained within this corridor to ensure safety. If the convex hulls fail to encompass the corresponding line segments, it may result in a discontinuity in the corridor, making it impossible to generate a continue trajectory within the corridor. Additionally, in whole-body planning [11, 3] that takes into account the shape of the robot, if the generated convex hulls fail to encompass the robot, it can lead to similar planning failure due to the absence of a solution space, for which we provide a example in Fig. 1(b). However, most existing algorithms prioritize optimizing the size of the region without adequately considering [2, 1] or ability to ensure manageability [9, 12].
@@
Figure 1: Illustration of the quality and manageability. The gray polytopes represent obstacles, and the green polytopes represent the generated obstacle-free convex polytopes.
@@
    Restrictive Inflation (RsI): inflating the ellipsoid and utilizing its contact planes tangent to the obstacles to generate a set of halfspaces, which separate a convex polytope from the obstacles.
@@
MVIE serves as a lower bound of the volume of the convex polytope, which monotonically expands during the iterative update, leading to a growing obstacle-free region that ensures the high quality of FIRI. The idea of monotonic updating the lower bound is inspired by Deits’ fundamental work IRIS [1], which, however, lacks manageability. Additionally, its considerable computational overhead in complex environments [1, 9] limits its real-time applications. In contrast, FIRI ensures manageability by restricting the halfspaces that composes the polytope to necessarily exclude obstacles yet contain the seed in RsI. In terms of efficiency, we design methods specifically for the two optimization-based modules involved in FIRI, leveraging their geometric properties, which results in a remarkable computational efficiency improvement, achieving a speedup of orders of magnitude compared to IRIS.
@@
MVIE, known as one of the most challenging extremal ellipsoid problems [17], is highly demanded in a number of applications [18, 19]. Most existing methods adopt semidefinite programming (SDP) formulations [20, 1] and employ various variants of interior point methods for solution [17, 21, 22, 23]. They face challenges in achieving satisfactory computational efficiency for massive constraint scenario, due to the need to handle with large-scale systems of equations in each iteration. To ease the computation overhead without sacrificing the solution quality, we present an equivalent formulation for MVIE in the form of second-order conic programming (SOCP). This brings about a noticeable boost in computational efficiency.
Moreover, especially in 2-D scenarios, we gain ultra computational efficiency by proposing an analytical method for MVIE. The proposed method exhibits a time complexity linear in the number of the hyperplanes of the input convex polytope, which is given for the first time to the best of our knowledge. As the dual problem of MVIE, 2-D Minimum Volume Enclosing Ellipse (MVEE) has long been equipped with linear-time analytical algorithms [24] based on its LP-type problem structure [25]. However, the corresponding approach for MVIE remains absent for several decades, due to its failure to satisfy the properties of LP-type problem and the lack of analytical solutions for basis computation [26]. In this paper, we address above challenges by designing an improved randomized algorithm with a careful problem reformulation and a bottom-up strategy inspired by the distance subalgorithm in the GJK algorithm [27]. As a result, the proposed specialized 2-D method achieves a substantial efficiency gain compared to other state-of-the-art methods [1, 23], surpassing them by several orders of magnitude.
Based on the above methods for RsI and MVIE, the proposed algorithm FIRI achieves efficiently generating high-quality convex polytope while ensuring manageability. To validate the performance of FIRI in these requirements, we conduct comprehensive comparisons between FIRI and various convex polytope generation algorithms [1, 2, 9]. The results provide compelling evidence that FIRI outperforms other approaches across the three aforementioned requirements. Moreover, we extensively perform real-world applications to showcase the applicability of FIRI, which involves a 2-D vehicle with non-holonomic constraint and a 3-D quadrotor, as well as point-mass model and whole-body model.
@@
    Restrictive Inflation (RsI) is introduced to ensure the manageability of the generated convex polytope. Based on its characteristic of few variables but rich constraints, a specialized and efficient solver is designed.
@@
    A novel method that formulates the MVIE problem into SOCP formulation is proposed, which avoids directly confronting the positive definite constraints and improves the computational efficiency.
@@
    Building upon the above methods, a novel convex polytope generation algorithm FIRI is proposed. Extensive experiments verify its superior comprehensive performance in terms of quality, efficiency, and manageability. High-performance implementation of FIRI will be open-sourced for the reference of the community.
@@ II-A Generating Free Convex Polytope
Deits et al. [1] propose Iterative Region Inflation by Semidefinite programming (IRIS) with the objective of generating the largest possible free convex polytope. IRIS involves inflating an ellipsoid, using it as a seed, to obtain a convex polytope formed by the intersection of contact planes. Then in the next iteration, the MVIE of the obtained convex polytope is selected as the new seed for inflation. However, this method requires solving semidefinite programming problems in each iteration to obtain the MVIE, which significantly hampers the computational efficiency. Furthermore, it does not guarantee that the generated convex hull always includes the initial seed point, thereby lacking manageability. Although IRIS is able to terminate iterations when the generated polytope does not contain the seed point, its essence does not lie in maximizing the volume while simultaneously constraining the seed to be inside. In contrast, FIRI approaches this problem from the perspective of constrained optimization, incorporating RsI that considers more general scenarios. Based on IRIS, Dai et al. [28] propose C-IRIS, which focuses on mapping collision in the task space to the configuration space based on kinematics and then generates certified safe convex hulls in the configuration space. This work is primarily applied to multi-joint robotic arms, which are different from the focus of this paper. To address the efficiency and local manageability deficiencies of IRIS, Liu et al. [2] propose the Regional Inflation by Line Search algorithm (RILS) which takes line segment as the seed input. RILS first generates a maximal ellipsoid that contains the segment yet excludes all obstacles and then inflates the ellipsoid to form a convex polytope with contact planes from the obstacles. The RILS algorithm demonstrates high computational speed. Nonetheless, since the ellipsoid always make the input line segment as major axis, it tends to produce conservative convex hull. Specifically for voxel maps, Gao et al. [8] develop Parallel Convex Cluster Inflation algorithm. The algorithm starts with a seed voxel that is unoccupied and then incrementally grows in layers along the coordinate axis using the visibility, which ensures the convexity of the voxel set throughout the growth process. Although it utilizes parallel computing to accelerate the computation, this algorithm can only achieve near real-time performance when the map resolution is low.
@@
On the other hand, Savin et al. [12] utilize the concept of space inversion. They apply a spherical polar mapping transformation to all the obstacle points with respect to the input seed point, which flips the feasible region around the seed point to the outside. Then, they compute the convex hull of the inverted obstacle point set and then transform it back to the original space to obtain the final convex hull. The limitation of this method lies in the strong nonlinearity of the spherical polar mapping, which means that any volume loss in the inverted space can result in significant gaps between the final convex polytope and the obstacles, leading to an overly conservative result. To address this limitation, Zhong et al. [9] propose using sphere flipping mapping as an alternative approach. Due to the properties of the hidden point remove operator inherent [29] in this mapping, this approach obtains the visible star-convex region around the seed point. They further devise a heuristic method to partition the star-convex region into the final convex polytope. However, the approach involves Quickhull [30] whose performance degrades when confronted with the points that are approximately distributed near the sphere after the mapping. In addition, heuristic partitioning can also lead to conservative result.
@@ II-B Maximum Volume Inscribed Ellipsoid (MVIE)
MVIE is also known as inner Löwner-John ellipsoid [31]. Nesterov et al. [22] utilize an interior-point algorithm with a specialized rescaling method on each iteration to achieve a polynomial time solution of MVIE, which is much faster than using ellipsoid algorithm [32]. Khachiyan et al. [33] transform the problem into a sequence of subproblems with only linear constraints, constructed by using the barrier method. This approach requires fewer computations compared to Nesterov’s method [22]. Then Anstreicher et al. [34] make improvements to both methods [22, 33] and demonstrate that computing an approximate analytic center of the polytope beforehand can reduce the complexity effectively. Zhang et al. [21] also provide a modification of Khachiyan’s approach [33] by replacing the inefficient primal barrier function method with a primal-dual interior-point method to solve the subproblems. Additionally, instead of dealing with a number of subproblems, they propose a novel primal-dual interior-point method free of matrix variables to solve MVIE directly. Through nonlinear transformations, they eliminate the positive-definite constraint on the coefficient matrix of the ellipsoid during the iteration process and provide a proof demonstrating that these nonlinear transformations preserve the uniqueness of the solution. Building upon similar idea of eliminating matrix variables, Nemirovskii [35] reformulates MVIE as a saddle-point problem based on the Lagrangian duality. Then they adopt a path-following method to obtain approximate saddle points according to the self-concordance theory proposed in Nesterov’s work [22]. However, these aforementioned interior-point methods face challenges when dealing with the scenario where the number of constraints is significantly larger than the dimension of the space. In these scenarios which are the focus of this paper, the large-scale system of linear equations that these methods need to solve at each iteration prevents them from computing MVIE within an acceptable time. To address the issue of inefficiency, Lin et al. [36] employ the fast proximal gradient method [37] to introduce a first-order optimization-based approach for MVIE. Although this approach significantly improves computational efficiency, it does not provide an exact solution since it approximates the non-differentiable indicator function by employing a differentiable one-sided Huber function to deal with the positive-definite constraint in MVIE.
MVIE is the most challenging problem among the extremal ellipsoid problems [17]. Other extremal ellipsoid problems, such as Minimum Volume Enclosing Ellipsoid (MVEE), can be transformed into MVIE with a linear reduction, which is irreversible. In addition to interior-point methods, analytical solutions of MVEE can be obtained using a linear-time randomized method in 2-D case [24]. However, for the more challenging MVIE, the corresponding linear complexity algorithm are absent for several decades. This gap will be addressed in this paper.

### v3

@@ Abstract
Convex polytopes have compact representations and exhibit convexity, which makes them suitable for abstracting obstacle-free spaces from various environments. Existing generation methods struggle with balancing high-quality output and efficiency. Moreover, another crucial requirement for convex polytopes to accurately contain certain seed point sets, such as a robot or a front-end path, is proposed in various tasks, which we refer to as manageability. In this paper, we propose Fast Iterative Regional Inflation (FIRI) to generate high-quality convex polytope while ensuring efficiency and manageability simultaneously. FIRI consists of two iteratively executed submodules: Restrictive Inflation (RsI) and Maximum Volume Inscribed Ellipsoid (MVIE) computation. By explicitly incorporating constraints that include the seed point set, RsI guarantees manageability. Meanwhile, iterative MVIE optimization ensures high-quality result through monotonic volume bound improvement. In terms of efficiency, we design methods tailored to the low-dimensional and multi-constrained nature of both modules, resulting in orders of magnitude improvement compared to generic solvers. Notably, in 2-D MVIE, we present the first linear-complexity analytical algorithm for maximum area inscribed ellipse, further enhancing the performance in 2-D cases. Extensive benchmarks conducted against state-of-the-art methods validate the superior performance of FIRI in terms of quality, manageability, and efficiency. Furthermore, various real-world applications showcase the generality and practicality of FIRI. The high-performance code of FIRI will be open-sourced.
@@ I Introduction
In robotics, a key task is to navigate without collisions, which involves frequent interactions with environments abundant in vast amounts of discrete obstacle information. For this interaction requirement, convex polytopes provide a compact, structured geometric abstraction of feasible space, alleviating the burden of collision avoidance [1, 2, 3]. Furthermore, this abstraction facilitates large-scale space storage and enables topological analysis, including the construction of road maps [4] and convex covers [5]. Additionally, convex polytopes construct convex linear safety constraints from non-convex obstacles, which benefits problem formulation and solution. In fact, in several applications, this approach even transforms the problem into convex optimization, leading to the attainment of global optima [6, 7].
@@
Although convex polytopes offer compact representation, generating satisfactory ones is far from a trivial task. First, a larger convex polytope extracts more information from the safety space, benefiting tasks that require spatial search within the safety space, such as trajectory planning. As illustrated in Fig. 1(a), larger convex polytopes facilitate a smoother trajectory. Therefore, we aim to maximize the polytope volume, which we define as its quality in this paper. Furthermore, maximizing computational efficiency enables applications in online high-speed tasks or with limited onboard resources. Despite extensive research [1, 8, 9, 2], balancing polytope quality and generation efficiency remains challenging. Existing approaches either generate high-quality regions but require substantial computational budget [1, 8], or achieve fast computation while yielding conservative results [9, 2].
@@
In addition to quality and efficiency, several applications impose a crucial requirement for the generated convex polytope to accurately encompass a specified set of points, to which we refer as the manageability. For example, in trajectory planning shown in Fig. 1(b), the corridor-based approach [2, 10] uses blue path segments as seeds to generate convex hulls, forming a safety corridor. If the convex hulls fail to encompass the corresponding line segments, it may result in a discontinuity in the corridor, making it impossible to generate a continue trajectory within the corridor. Additionally, in whole-body planning [11, 3], insufficient coverage of the robot’s shape can lead to similar planning failure due to the absence of feasible solution space, as illustrated in Fig.1(b). However, most existing algorithms prioritize optimizing the size of the region without adequately considering [2, 1] or ability to ensure manageability [9, 12].
@@
Figure 1: Illustration of quality and manageability. The gray polytopes represent obstacles, and the green polytopes represent the generated free polytopes.
@@
    Restrictive Inflation (RsI): inflating the ellipsoid and using its contact planes tangent to obstacles to generate a set of halfspaces containing the seed, which separate a convex polytope from obstacles.
@@
MVIE serves as a lower bound of the volume of the convex polytope, which monotonically expands during the iterative update, leading to a growing obstacle-free region that ensures the high quality of FIRI. The idea of monotonic updating the lower bound is inspired by Deits’ fundamental work IRIS [1], which, however, lacks manageability. Additionally, its considerable computational overhead in complex environments [1, 9] limits its real-time applications. In contrast, FIRI significantly improves upon IRIS in terms of manageability and computational efficiency. For manageability, FIRI uses RsI to generate halfspaces that compose the polytope to necessarily exclude obstacles while containing the seed. For efficiency, we design specialized methods for FIRI’s two optimization modules based on their geometric properties, resulting in a remarkable computational efficiency improvement with a speedup of orders of magnitude compared to IRIS.
@@
MVIE, known as one of the most challenging extremal ellipsoid problems [17], is highly demanded in a number of applications [18, 19]. Most existing methods adopt semidefinite programming (SDP) formulations [20, 1] with interior point method variants for solution [17, 21, 22, 23]. However, they struggle with computational efficiency when handling massive constraint due to large equation systems required in each iteration. To reduce computational overhead without sacrificing solution quality, we present an equivalent second-order conic programming (SOCP) formulation for MVIE, bringing about a noticeable boost in computational efficiency.
Moreover, especially in 2-D scenarios, we gain ultra computational efficiency by proposing an analytical method for MVIE. The method exhibits a time complexity linear in the number of the hyperplanes of the input convex polytope, which is given for the first time to the best of our knowledge. As the dual problem of MVIE, Minimum Volume Enclosing Ellipse (MVEE) has long been equipped with linear-time analytical algorithms [24] based on its LP-type problem structure [25]. However, the corresponding approach for MVIE remained absent for decades due to failing LP-type problem properties and lacking analytical solutions for basis computation[26]. We address these challenges through an improved randomized algorithm with problem reformulation and a bottom-up strategy inspired by GJK algorithm’s distance subalgorithm [27]. Consequently, the specialized 2-D method achieves a substantial efficiency gain compared to other state-of-the-art methods [1, 23], surpassing them by several orders of magnitude.
Building upon the above methods for RsI and MVIE, FIRI generates high-quality convex polytopes with manageability while maintaining high efficiency. To evaluate the performance of FIRI in these three requirements, we conduct comprehensive comparisons with various polytope generation algorithms [1, 2, 9]. The results provide compelling evidence that FIRI outperforms other approaches across all three requirements. Moreover, we extensively perform real-world applications to showcase the applicability of FIRI, which involves a 2-D vehicle with non-holonomic constraint and a 3-D quadrotor, as well as point-mass and whole-body model.
@@
    RsI is introduced to ensure the manageability of the generated polytope. For its few-variable but rich-constraint nature, a specialized and efficient solver is designed.
@@
    A novel SOCP formulation for MVIE is proposed, avoiding the positive definite constraints and improving the computational efficiency.
@@
    Building upon the above methods, a novel convex polytope generation algorithm FIRI is proposed. Extensive experiments verify its superior comprehensive performance in terms of quality, efficiency, and manageability.
@@ II-A Generating Free Convex Polytope
Deits et al. [1] propose Iterative Region Inflation by Semidefinite programming (IRIS) with the objective of generating the largest possible free convex polytope. IRIS involves inflating an ellipsoid, using it as a seed, to obtain a convex polytope formed by the intersection of contact planes. Then in the next iteration, the MVIE of the obtained convex polytope is selected as the new seed for inflation. However, this method requires solving semidefinite programming problems in each iteration to obtain the MVIE, which significantly hampers the computational efficiency. Furthermore, the hyperplane computation of IRIS cannot directly accommodate additional seed containment constraint, lacking manageability. In contrast, our RsI for FIRI unifies the forms of obstacle exclusion and seed containment constraints, enabling simpler and more efficient solutions. Wu et al. [28] use an IRIS-like method to obtain convex polytope through ellipsoid-based hyperplanes. They propose an ADMM-based approach that iteratively refines polytopes and optimizes trajectories to achieve shorter and better trajectory. Based on IRIS, Dai et al. [29] propose C-IRIS, which focuses on mapping collision in the task space to the configuration space based on kinematics and then generates certified safe convex hulls in the configuration space. Similarly, Mark et al.[30] propose IRIS-NP, extending IRIS to configuration space via nonlinear programming. Subsequently, building upon IRIS-NP, Werner et al.[31] significantly accelerate polytope generation by sampling nearby configuration-space obstacles. The above IRIS derivatives primarily target multi-joint robotic arms in configuration space, differing from our focus. To address the efficiency and local manageability deficiencies of IRIS, Liu et al. [2] propose the Regional Inflation by Line Search algorithm (RILS) which takes line segment as the seed input. RILS first generates a maximal ellipsoid that contains the segment yet excludes all obstacles. Then RILS inflates the ellipsoid to form a convex polytope with contact planes from the obstacles, which fundamentally aligns with the inflation step in a single iteration of IRIS. RILS shows high computational speed. Yet using input line as major axis of the ellipsoid, it tends to produce conservative convex hull. Specifically for voxel maps, Gao et al. [8] develop Parallel Convex Cluster Inflation algorithm. Starting from an unoccupied seed voxel, it grows incrementally in layers along coordinate axis using visibility, maintaining voxel set convexity. Despite parallel computing acceleration, this algorithm reaches near real-time only at low map resolution.
@@
On the other hand, Savin et al. [12] utilize the concept of space inversion. Using spherical polar mapping from the input seed point, they flip all the obstacle points outside. After computing convex hull of the inverted obstacle points, they transform it back to the original space for the final hull. The strong nonlinearity of spherical polar mapping limits this method, as volume loss in inverted space creates gaps between final polytope and obstacles, yielding conservative results. To address this limitation, Zhong et al. [9] adopt sphere flipping mapping. Due to the properties of the hidden point remove operator inherent [32] in this mapping, this approach obtains the visible star-convex region around the seed. They then partition this star-convex region into final convex polytope through a heuristic method. Yet this approach uses Quickhull [33], which degrades with points distributed near the sphere after mapping. In addition, the heuristic partitioning yields conservative result.
@@ II-B Maximum Volume Inscribed Ellipsoid (MVIE)
MVIE is also known as inner Löwner-John ellipsoid [34]. Nesterov et al. [22] utilize an interior-point algorithm with a specialized rescaling method on each iteration to achieve a polynomial time solution of MVIE, surpassing ellipsoid algorithm [35]. Khachiyan et al. [36] transform the problem into a sequence of subproblems with only linear constraints, constructed by using the barrier method. This approach requires fewer computations compared to Nesterov’s method [22]. Then Anstreicher et al. [37] make improvements to both methods [22, 36] and demonstrate that computing an approximate analytic center of the polytope beforehand can reduce the complexity effectively. Zhang et al. [21] also provide a modification of Khachiyan’s approach [36] by replacing the inefficient primal barrier function method with a primal-dual interior-point method to solve the subproblems. Additionally, instead of dealing with a number of subproblems, they propose a novel primal-dual interior-point method free of matrix variables to solve MVIE directly. Through nonlinear transformations, they eliminate the positive-definite constraint on the coefficient matrix of the ellipsoid during the iteration process and provide a proof demonstrating that these nonlinear transformations preserve the uniqueness of the solution. Building upon similar idea of eliminating matrix variables, Nemirovskii [38] reformulates MVIE as a saddle-point problem using Lagrangian duality. They then use path-following method for finding approximate saddle points based on Nesterov’s self-concordance theory [22]. However, these interior-point methods struggle with scenarios having constraints far exceeding space dimension. In these scenarios which are the focus of this paper, the large-scale system of linear equations that these methods need to solve at each iteration prevents them from computing MVIE within an acceptable time. To address inefficiency, Lin et al. [39] employ the fast proximal gradient method [40] to introduce a first-order optimization-based approach for MVIE. Although this approach significantly improves efficiency, it lacks exact solutions due to approximating non-differentiable indicator with an one-sided Huber function for positive-definite constraints.
MVIE is the most challenging problem among the extremal ellipsoid problems [17]. Other extremal ellipsoid problems, such as Minimum Volume Enclosing Ellipsoid (MVEE), can be transformed into MVIE with a linear reduction, which is irreversible. Beyond interior-point methods, analytical solutions of MVEE can be obtained using a linear-time randomized method in 2-D case [24]. However, for the more challenging MVIE, the corresponding linear complexity algorithm are absent for several decades. This paper will address this gap.